3 large popcorns and 2 small drinks at the movie theatre cost $21.00. 2 large popcorn and 4 small drinks costs $22.00. What is the cost of 1 small drink?
Answer by hamsanash1981@gmail.com(151)

let popcorns be x and small drinks be y
then, 3x+2y = 21 ---1
2x + 4y = 22 ---2
Solving equation 1 and 2 we have
6x + 4y = 42
6x +12y = 66
============
8y = 24
y = 3
therefore the small drink cost $3.00